OpenStudy (anonymous):

Can anybody give me a general definition of a Maclaurin series and solve the exampleof f(x)=sin(2x)

OpenStudy (tkhunny):

Can you do \(h(x) = \sin(x)\)?

OpenStudy (anonymous):

No I don't understand the series at all

OpenStudy (tkhunny):

Well, that's not very encouraging. We just need a lot of derivatives. Show me the first four derivatives of your f(x) and we can talk.

OpenStudy (anonymous):

Cos(2x)(2)

OpenStudy (anonymous):

No?

OpenStudy (tkhunny):

That's one. Three more.

OpenStudy (anonymous):

So is that all just the repetition of the derivative watt is the propose/final product

OpenStudy (anonymous):

–4sin(2x) –8cos(2x) 16Sin(2x)

OpenStudy (tkhunny):

That's about it. Now you just have to build it. \(\dfrac{\sin(2\cdot 0)}{0!}x^{0} + \dfrac{2\cdot \cos(2\cdot 0)}{1!}x^{1} + \dfrac{-4\cdot \sin(2\cdot 0)}{2!}x^{2} + \dfrac{-8\cdot \cos(2\cdot 0)}{3!}x^{3} + ...\) Eventually, you should notice two things: 1) All the terms with sin(0) simply drop out. 2) The series is alternating.

OpenStudy (anonymous):

So the steps to constructing this is to fill in x with zero then put each derivative over increasing terrorists?

OpenStudy (anonymous):

Darn auto correct mental factorial

OpenStudy (anonymous):

Ment factorial

OpenStudy (tkhunny):

You evaluate each derivative at the desired point. In this case, x = 0 because we said "Maclaurin" in the Problem Statement.

OpenStudy (anonymous):

Ok I think I have a decent idea thank you very much for your time and knowledge